# inkscapelatex=false

### v1

@@ 4 GCN-DevLSTM Network for SAR
We illustrate the pipeline of our method in Figure 2(a). As shown in Figure 2(a), we treat the GCN and the DevLSTM module as an integrated block. GCN module focuses on capturing spatial correlations among vertices, while DevLSTM module is designed to capture temporal dynamics of the temporal graph across video frames. Detailed explanations of the GCN and DevLSTM modules will be provided in the subsequent sections. Our network is designed to capture discriminative spatial-temporal features with $N$ such blocks, where $N$ is set to 9 in this paper. Additionally, within each block, we build a residual path that establishes a shortcut connection with the input, which can mitigate issues related to gradient vanishing and explosion, enhancing the training stability of the model.
@@ 4.1 GCN Module
In the GCN module, we basically follow the structure of CTR-GC Chen et al. (2021). However, we depart from the use of three initial adjacency matrices representing self, inward, and outward correlations. Instead, we enhance the model by incorporating various hops of information to construct three adjacency matrices denoted as $A_{0}$, $A_{1}$, and $A_{2}$.
@@
$A_{0}$ remains an identical (self) matrix. For $A_{1}$, the elements $a_{ij}$ are set to 1 if vertex i and vertex j are physically connected in the skeleton graph; otherwise $a_{ij}$ in $A_{1}$ is set to 0. $A_{1}$ captures the direct connections in the graph. To enrich spatial information, we introduce $A_{2}=A_{1}^{2}$, enabling a vertex to connect with its neighbour’s neighbour. $A_{2}$ allows us to extend our understanding beyond direct connections to include relationships involving two hops in the graph.
Following CTR-GC, we utilise a network to dynamically learn an adjacency matrix for each channel, adding it with $A_{k}$ for $k\in\{0,1,2\}$. This process can be expressed as:
$A_{k}^{{}^{\prime}}(X)=A_{k}+\mathcal{F_{\theta}}(X)$ (5)
where $A_{k}^{{}^{\prime}}$ represents the updated adjacency matrix, $A_{k}$ is the initial adjacency matrix for the $k$-th hop, and $\mathcal{F_{\theta}}(X)\in\mathbb{R}^{M\times M\times C}$ dynamically captures the spatial correlations of joints from the network, conditioned on the input data $X\in\mathbb{R}^{T\times M\times d}$, considering each channel of the output channels $C$. An overview of the CTR-GC is given in supplementary while details about the network of $\mathcal{F_{\theta}}(X)$ can be found in Chen et al. (2021).
After obtaining the input-conditioned adjacency matrix, the graph convolution is further performed as
$Z_{t,c}=A_{k}^{{}^{\prime}}(X)_{c}\cdot(X_{t}\cdot W_{s})_{c}$ (6)
where $Z\in\mathbb{R}^{T\times M\times C}$ is the output of GCN with the feature dimension $C$, and $W_{s}\in\mathbb{R}^{d\times C}$ denotes a feature transformation filter. We sum the outputs from three CTR-GC modules, each employing a distinct adjacency matrix. Subsequently, we apply batch normalisation and ReLU activation function to the summation result before sending it to the DevLSTM module.
4.2 DevLSTM Module
The core of our DevLSTM module is path development layer, where we employ a sliding-window-based strategy for processing time sequence data. Specifically, for each joint $V^{i}$, given time sequence data $Y\in\mathbb{R}^{T\times d}$, window (kernel) size $K$, dilation $D$, padding size $P$ and stride $S$, for $10\leq i\leq\frac{T-K-(K-1)(D-1)+2P}{S}-$ 1, we define each partition $Z_{i,D}\in Z$ as:
$Z_{i,D}=[Y_{i\times S+p},Y_{i\times S+p+D},\dots,Y_{i\times S+p+(K-1)\cdot D}]$ (7)
where $p$ is the offset to account for padding. To guarantee that we will not lose any information related to changes in the path, it is crucial to ensure a minimum one-point overlap between two consecutive time partitions.
We give an example in Figure 2(c) to clearly illustrate how we partition the given time sequence and process these partitions with path development layer. In the given example, the time sequence data $Y$ from a joint across time is segmented in a sliding-window way with the setting $K=4$, $D=1$, $S=2$, and at least one-point overlap is guaranteed. One shared path development layer then maps these partitions to the Lie group using Eq. 4 in a recurrent fashion. For each partition, we only retain the path development layer’s state at the final time point, ignoring the intermediate stages and the state will be further flattened from $m\times m$ to $1\times m^{2}$. We concatenate all flattened states from these partitions along the temporal dimension, thereby producing a new time sequence with the dimension $\mathbb{R}^{T{{}^{\prime}}\times m^{2}}$.
Inspired by the design of multi-scale tcn Liu et al. (2020), we also design a multi-scale network to capture temporal dynamics. We utilise $D=1$ and $D=2$ respectively to enable the path development layer to capture features at different time scales, enhancing the module’s ability to discern temporal patterns across different scales in the input sequence. Additionally, it’s worth noting that zero padding, a commonly employed technique in CNNs, may not be the most suitable choice for the path development layer from experimental results. In our approach, we opt for replicate padding.
Due to the translation invariance of the path development, we append the location of the starting points ($SP=Z_{D,0}$) to the output of the path development layer along the feature dimension over each sub-time interval. In formula, it can be expressed as
$C_{i}=Concat(Dev_{\theta_{1}}(Z_{i,D=1}),Dev_{\theta_{2}}(Z_{i,D=2}),SP_{i})$ (8)
where $C_{i}$ is the concatenation result at partition $i$, $Dev_{\theta_{1}}$ and $Dev_{\theta_{2}}$ are two distinct path development layers regarding to $D=1$ and $D=2$, respectively, and $SP_{i}$ is the first element of partition $i$. This step is crucial for uniquely determining a path, enhancing the effectiveness of the path development layer in capturing relevant temporal features.
@@ 4.2 DevLSTM Module
Our method(J) 90.8 95.7 86.2 87.7
Our method(B) 91.1 95.2 87.0 88.6
Our method(D) 91.2 95.3 87.1 88.7
Our method(J+B) 92.4 96.4 89.0 90.4
Our method(J+B+JM+BM) 92.6 96.6 89.3 90.7
Our method(J+B+D+JM+BM) 92.9 96.6 89.7 91.2
The concatenated results $C=\{C_{0},C_{1},...,C_{N}\}$ where $N=\frac{T-K-(K-1)(D-1)+2P}{S}-1$, from two path development layers, along with their corresponding starting points, are subsequently fed into an LSTM network whose unit is crafted to selectively learn and store information over long sequences, enabling it to grasp long-term dependencies in data. There won’t be any changes to the temporal dimension before and after LSTM because all of LSTM’s hidden states at each time step will be utilised.
MaxPool and a convolutional layer with $5\times 1$ kernel size have consistently demonstrated superior performance within the temporal module of other GCN-based studies. Introducing both MaxPool and the $5\times 1$ convolutional operation to our DevLSTM temporal module has the potential to enhance model performance even further. We can reasonably design kernel size, padding, stride and dilation to ensure the output size of Maxpool, LSTM and $5\times 1$ convolution remain the same. As a result, we can sum up their outcomes to achieve improved overall performance.
5 Numerical Experiments
To validate the effectiveness of the proposed GCN-DevLSTM network, we carry out numerical experiments on the following popular benchmarking datasets for skeleton-based action and gesture recognition. (1) NTU RGB+D 60 (NTU-60) Shahroudy et al. (2016) consists of 56,880 action samples and 60 action classes, filmed by 40 people using three Kinect cameras in different camera views. 3D skeletal data in the dataset have different sequence lengths, with each owning 25 joints. (2) NTU RGB+D 120 (NTU-120) Liu et al. (2019b) extends upon NTU RGB+D 60, consisting of 114,480 samples and 120 action classes, filmed by 106 people in three camera views. We consider the Cross-subject (X-Sub) and Cross-view (X-View) evaluation benchmarks for both datasets as proposed by Shahroudy et al. (2016). (3) Chalearn2013 Escalera et al. (2013) is a public dataset for gesture recognition, consisting of 11,116 skeleton examples and 20 gesture categories, filmed by 27 people. We use only skeleton data from these datasets for our experiments.
Experiments are conducted using a single Nvidia A6000 graphics card, where a batch of 16 examples is fed to the network. Stochastic gradient descent optimiser Bottou (2012) is employed, with a Nesterov $\text{momentum}=0.9$ and a weight decay $=0.0003$. The training epoch is set to 65 and we apply the warm-up strategy He et al. (2016) to the first 5 epochs. The initial learning rate is 0.02 and a step learning rate decay is applied with a factor of 0.1 at epoch 35 and 55, respectively. Data processing on NTU datasets mainly follows the procedure in Chen et al. (2021), with the length of all input data sequences resizing to 64. We use $\mathfrak{se}$ Lie algebra in this paper, with a corresponding matrix size of $m=10$. A detailed network architecture can be found in the supplementary.
5.1 Dual graph
[…]
@@ 5.2 Comparison with state-of-the-art methods
We compare our approach with other state-of-the-art methods on both NTU-60 and NTU-120 datasets, as shown in Table 1. Since our GCN module is following CTR-GCN Chen et al. (2021) with the major difference lying in the temporal module, we show more results on CTR-GCN for a more comprehensive comparison. As we can see all of our results, no matter whether using only joint or using ensembled streams combining joint and bone, outperform CTR-GCN with a large margin on all benchmarks. Our method’s limited performance gain with extra joint or bone motion, compared to CTR-GCN, may stem from our path development layer exploiting position differences between consecutive time points. As a result, streams using joint and bone motion provide identical information at the beginning of the network, limiting their help. Our model, which integrates five data streams, achieves the best result compared to other state-of-the-art approaches across three benchmarks, with the exception of a close performance on the NTU60-Xview benchmark.
Table 2 shows the comparison results between our result and other state-of-the-art methods on Chalearn2013 dataset. We only employ the joint stream for fair comparison as other approaches only use the joint representation on Chalearn2013 dataset. It is evident that our method yields the best result again on the gesture dataset, further proving its superiority.
Table 2: Experimental results on Chalearn2013 dataset.
@@
ST-PSM+L-PSM [^1] Other methods all use data augmentation except the one from Cheng et al. (2023). However, their code is not publicly available, resulting in no way to retrain their approach using augmented data. Cheng et al. (2023) 94.18
@@ 5.4 Comparison with MS-TCN
5.5 Ablation studies
We further investigate the importance of each component in our temporal module. As evident from Table 4, the path development layer emerges as the primary contributor to the overall performance improvement since removing it causes the largest performance degradation, with other components also contributing to enhanced performance.
Table 4: Ablation studies of DevLSTM module on NTU-60 dataset.
@@ 5.6 Robustness Analysis
We compare the robustness of our full model (GCN-DevLSTM), CTRGCN Chen et al. (2021) and our model without path development layer (GCN-LSTM) when missing frames happen. To this end, we create a new test dataset by randomly dropping a certain number of frames out of 64 frames and validate the model performance on this test dataset. Figure 3 illustrates that our full model with the path development layer consistently outperforms the other methods in terms of accuracy, regardless of the proportion of frames missing, proving its robustness to irregular sampling. More results are available in the supplementary material.
Figure 3: Robustness analysis against the dropping frames on NTU60-Xsub benchmark.
6 Conclusion
In summary, we propose a novel and general GCN-DevLSTM network for analysing temporal graph data by leveraging the path development layer. This approach not only offers flexibility when combined with various GCN modules but also significantly enhances performance by effectively capturing the temporal dependencies within data. The effectiveness of the proposed method is validated by the state-of-the-art results on both small gesture data and large action datasets in SAR. Moreover, it demonstrates superior robustness to missing data and variable frame rates.
Acknowledgements. LJ, WY and HN are supported by the EPSRC [grant number EP/S026347/1]. WY and HL are also supported by The Alan Turing Institute under the EPSRC grant EP/N510129/1.

### v2

@@ 4 GCN-DevLSTM Network for SAR
In Figure 2(a), we introduce the proposed GCN-DevLSTM Network, a novel framework for SAR tasks that integrates the GCN and DevLSTM modules into a single block. The GCN module is designed to detect spatial correlations among vertices, while the DevLSTM module handles the temporal dynamics across video frames. Further details on these modules are discussed in Sections 4.1 and 4.2, respectively. Our network is designed to capture discriminative spatial-temporal features with $N$ such blocks, where $N=9$ in this paper. Additionally, within each block, we build a residual path that establishes a shortcut connection with the input to mitigate gradient vanishing/ explosion issues, enhancing the training stability.
@@ 4.1 GCN Module
In the GCN module, we basically follow the network of CTR-GC Chen et al. (2021). However, we depart from the use of three initial adjacency matrices representing self, inward, and outward correlations. Instead, we propose a Hops-CTRGC to enhance the model by incorporating various hops of information to construct three adjacency matrices denoted as $A_{0}$, $A_{1}$, and $A_{2}$.
@@
$A_{0}$ remains an identical (self) matrix. For $A_{1}$, the elements $a_{ij}$ are set to 1 if vertex i and vertex j are physically connected in the skeleton graph; otherwise, $a_{ij}$ in $A_{1}$ is set to 0. $A_{1}$ captures the direct connections in the graph. To enrich spatial information, we introduce $A_{2}=A_{1}^{2}$, enabling a vertex to connect with its neighbour’s neighbour. $A_{2}$ allows us to extend our understanding beyond direct connections to include relationships involving two hops in the graph. Since GCN module is not our main focus in this paper, readers can find more details about the network of CTR-GC in Chen et al. (2021) or in our supplementary material 7.5.
4.2 G-DevLSTM Module
The proposed DevLSTM module, illustrated in Figure 2(b) is mainly composed of the CNN-based branch and the G-development-based branch. The output of DevLSTM module is given by applying GeLU activation function to the summation of the output of both branches.
G-Development branch. Inspired by the design of multi-scale TCN Liu et al. (2020), we also design a multi-scale network to capture temporal dynamics. We utilize two G-dev layers with the dilation parameter $D=1$ and $D=2$ to enhance its ability to capture temporal patterns across different scales in the input sequence.
The path development combined with the starting point can determine the path uniquely. Hence, we append the location of the starting points to the output of the path development layer along the feature dimension over each sub-time interval. The resulting features are then fed into an LSTM network, which is able to effectively aggregate the local development summary and learn deep features to extract the long-term dependencies in time series data. We use the sequential output of LSTM, which keeps the time dimension the same as the input.
CNN-based branch. MaxPool and a convolutional layer with $5\times 1$ kernel size have consistently demonstrated superior performance within the temporal module of other GCN-based studies. Introducing both MaxPool and the $5\times 1$ convolutional operation to our DevLSTM temporal module has the potential to enhance model performance even further. We can reasonably design kernel size, padding, stride and dilation to ensure the output size of Maxpool, LSTM and $5\times 1$ convolution remain the same.
5 Numerical Experiments
@@
Our methods Joint Bone Dual Joint motion Bone motion
90.8 95.7 86.2 87.7
91.1 95.2 87.0 88.6
91.2 95.3 87.1 88.7
92.4 96.4 89.0 90.4
92.6 96.6 89.3 90.7
92.9 96.6 89.7 91.2
To validate the effectiveness of the proposed approach, we carry out numerical experiments on the following popular benchmarking datasets for skeleton-based action and gesture recognition. These datasets are chosen for their diversity in data size and task complexity, providing a comprehensive evaluation of the model’s performance. (1) Chalearn2013 Escalera et al. (2013) is a public dataset for gesture recognition, consisting of 11,116 skeleton examples and 20 gesture categories, filmed by 27 people. Each frame has 20 joints. (2) NTU RGB+D 60 (NTU-60) Shahroudy et al. (2016) consists of 56,880 action samples and 60 action classes, filmed by 40 people using three Kinect cameras in different camera views. 3D skeletal data in the dataset have different sequence lengths, with each owning 25 joints. (3) NTU RGB+D 120 (NTU-120) Liu et al. (2019b) extends upon NTU RGB+D 60, consisting of 114,480 samples and 120 action classes, filmed by 106 people in three camera views. We consider the Cross-subject (X-Sub) and Cross-view (X-View) evaluation benchmarks for both NTU-60 and NTU-120 as proposed by Shahroudy et al. (2016). We use only skeleton data from these datasets for our experiments.
Experiments are conducted using a single Nvidia A6000 graphics card, where a batch of 16 examples is fed to the network. Stochastic gradient descent optimiser Bottou (2012) is employed, with a Nesterov $\text{momentum}=0.9$ and a weight decay $=0.0003$. The training epoch is set to 65 and we apply the warm-up strategy He et al. (2016) to the first 5 epochs. The initial learning rate is 0.02 and a step learning rate decay is applied with a factor of 0.1 at epoch 35 and 55, respectively. Data processing on NTU datasets mainly follows the procedure in Chen et al. (2021), with the length of all input data sequences resizing to 64. We use $\mathfrak{se}$ Lie algebra in this paper, with a corresponding matrix size of $m=10$. A detailed network architecture can be found in the supplementary. The replicate padding is used in our experiments.
5.1 Fusion results from different data streams
Most of the recent approaches that we are set to compare combine predicted scores from different input data streams for enhanced performance. The most commonly used data streams are joint (J), bone (B), joint motion (JM) and bone motion (BM) where details can be found in Shi et al. (2019, 2020). Apart from the above-mentioned data streams, we introduce a novel stream: the dual graph (D). In a dual graph, vertices represent bones, and edges establish connections between these bones. More details about dual graph can be found at Supplementary 7.1.
@@ 5.2 Comparison with state-of-the-art methods
Table 2: Results on Chalearn2013 dataset.
@@
ST-PSM+L-PSM Cheng et al. (2023) 94.18
@@
We compare our approach with other state-of-the-art methods on both NTU-60 and NTU-120 datasets, as shown in Table 1. Since our GCN module is following CTR-GCN Chen et al. (2021) with the major difference lying in the temporal module, we show more results on CTR-GCN for a more comprehensive comparison. As we can see all of our results, no matter whether using only joint or using ensembled streams combining joint and bone, outperform CTR-GCN with a large margin on all benchmarks. Our method’s limited performance gain with extra joint or bone motion, compared to CTR-GCN, may stem from our path development layer exploiting position differences between consecutive time points. As a result, streams using joint and bone motion provide identical information at the beginning of the network, limiting their help. Our model, integrating five data streams, achieves the best result compared to other state-of-the-art approaches across three out of four benchmarks.
Figure 3: Robustness analysis on NTU60 X-sub benchmark.
Table 2 shows the comparison results between our result and other state-of-the-art methods on Chalearn2013 dataset. We only employ the joint stream for fair comparison as other approaches only use the joint representation on Chalearn2013 dataset. It is evident that our model yields the best result on the small-scale gesture dataset, proving its effectiveness on the small dataset. It is important to highlight that our network structure remains the same on all datasets, demonstrating its strong adaptability for different dataset sizes.
@@ 5.4 Comparison with MS-TCN
Table 4: Ablation studies of DevLSTM module on NTU-60 dataset. w/o means without.
@@
5.5 Ablation studies
We further investigate the importance of residual path design to connect the input and output of each block and each component in our temporal module. As evident from Table 4, residual path plays an important role to enhance the model performance since it can help alleviate gradient vanishing/explosion in the deep neural network. For temporal components in our DevLSTM module, the path development layer emerges as the primary contributor to the overall performance improvement since removing it causes the largest performance degradation, with other components also contributing to enhanced performance.
@@ 5.6 Robustness Analysis
We comprehensively compare the robustness of our full model (GCN-DevLSTM), CTRGCN Chen et al. (2021) and our model without path development layer (GCN-LSTM) in terms of missing frames, missing joints and variable sequence length. To this end, we firstly create a new test dataset by randomly dropping or inserting a certain number of frames out of 64 frames to test the robustness of missing frames and variable length. Figure 3a illustrates that our full model with the path development layer consistently outperforms the other methods in terms of accuracy, regardless of the proportion of frames missing or inserting, proving its robustness to irregular sampling and variable sequence length.
Additionally, to validate the robustness of missing joints due to occlusions in real life, we create another dataset by randomly dropping a certain number of joints in each frame. To fill in these missing joints, two techniques are adopted: zero-filling (in Figure 3b) and linear interpolation (in Figure 3c). Both figures show that our model with G-Dev layer, is consistently more robust against missing joints than others. More details are in the supplementary material.
6 Conclusion and future work
Conclusion. In summary, we propose a novel and general GCN-DevLSTM network for analyzing temporal graph data by leveraging the path development layer. This approach not only offers flexibility when combined with various GCN modules but also significantly enhances performance by effectively capturing the temporal dependencies within data. The effectiveness of the proposed method is validated by the state-of-the-art results on both small gesture data and large action datasets in SAR. Moreover, it demonstrates superior robustness to missing data and variable frame rates. Limitation. The current G-Dev Sequence layer relies on other GCNs to extract spatial information, which limits its standalone utility. Future enhancements could include integrating the features of adjacent nodes directly into the input of the path development layer to better capture spatio-temporal dependencies. Broader impact Our proposed G-Dev Sequence layer provides a generic model for temporal graph data, opening up possibilities for its applications across various fields, including urban planning, social network analysis and others.
Acknowledgements. LJ, WY and HN are supported by the EPSRC [grant number EP/S026347/1]. WY and HN are also supported by The Alan Turing Institute under the EPSRC grant EP/N510129/1.